# Long-time behavior for discretization schemes of Fokker–Planck equations via couplings

### v1

@@ Abstract.
The authors acknowledge partial support from the Austrian Science Fund (FWF), grants P33010 and F65. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme, ERC Advanced Grant no. 101018153.
@@ 1.1. The setting
[…]
@@
where $G_{i}$ is the set of all moves ±$\pm_{j}$ to the neighboring states $i\mapsto i\pm he_{j}$ (with the grid size $h>0$) and $c(i,\gamma)$ denotes the rate that a jump happens from $i$ in the direction $\gamma$,
@@ 1.2. Main results and key ideas
Condition (i) is satisfied for the power-law entropies $\phi_{\alpha}(x)=x^{\alpha}$ with $1<\alpha\leq 2$ and the Boltzmann entropy density $\phi_{\alpha}(x)=x\log x$ with $\alpha=1$. For this family of functions, the exponential decay is closely related to convex Sobolev inequalities, namely the modified logarithmic Sobolev inequality for $\alpha=1$, the Poincaré inequality for $\alpha=2$, and Beckner-type inequalities for $21<\alpha<$ 2. The convexity of $\phi$ is natural in this context, while the convexity of $\Phi$ is needed in Conforti’s coupling approach; see Theorem 15 below. The strong convexity of the potential in Condition (ii) is required to conclude exponential decay. Condition (iii) is satisfied for additive potentials being of the form $V(x)=\sum_{j=1}^{d}V_{j}(x_{j})$ with $x=(x_{1},\ldots,x_{d})\in{\mathbb{R}}^{d}$; see Remark 5.
Our main results include the quantitative decay rates for the continuous-time Markov chain and its discrete-time version (obtained from an explicit Euler scheme) in the relative $\phi$-entropy
@@
(6) $\mathcal{H}^{\phi}(S_{t}f|m_{h})\leq e^{-\kappa_{\phi}t}\mathcal{H}^{\phi}(f|m_{h})$ for $\quad\mbox{for }f\geq 0,\ t\geq 0,$
(7) $\mathcal{W}_{1}(\nu p_{t},\eta p_{t})\leq\sqrt{d}e^{-\kappa_{1}t}\mathcal{W}_{1}(\nu,\eta)$ for $\quad\mbox{for }t\geq 0,$
@@
Therefore, the second idea is to use the coupling method of Conforti [14]. Roughly speaking, the idea is to compare two probability measures $\nu$ and $\eta$ and to construct a joint probability spaces with marginals $\nu$ and $\eta$. Conforti proved (6) for Markov chains with transition rates (5) but with a different value of $\kappa_{\phi}$. We recalculate $\kappa_{\phi}$ and interpret its value in the context of finite-volume schemes with strongly convex potentials. In the proof, we use a synchronous coupling for the contraction rates with maximal probability. When a synchronous move is not possible, Otherwise, we couple the rates of the neighboring states such that the states after a jump are identical with maximal probability or are at least neighboring states.
Inequality (7) is proved by using the path coupling method of [9]. It holds for two continuous-time Markov chains $Y_{t}^{1}$ and $Y_{t}^{2}$ that
@@
We prove similar results as (6)–(7) for discrete-time Markov chains associated to an explicit Euler scheme for the Fokker–Planck equation. Here, the idea is to apply the discrete Bakry–Emery method of [22], which is based on an estimate between the Fisher information $\mathcal{E}(\phi^{\prime}(f),f)$ and the entropy production $-({\mathrm{d}}\mathcal{H}^{\phi}/{\mathrm{d}}t)(S_{t}f|m_{h})$ (which are the same in the continuous setting). We do not obtain contraction but exponential decay with a prefactor that depends on the initial datum $f$.
@@ 1.3. Comparison to the literature
Mielke investigated in [26] geodesic convexity properties of nonlocal transportation distances on probability spaces such that continuous-time Markov chains can be formulated as gradient flows. It is known that the geodesic convexity property implies exponential decay [2]. This idea was extended to nonlinear Fokker–Planck equations in [12].
@@
Conforti established in [14] a new probabilistic approach to prove convex Sobolev inequalities of the type (8) and to quantify the exponential decay in terms of the entropy for continuous-time Markov chains. His approach uses coupling rates to treat the second time derivative of the entropy, thus avoiding the use of discrete Bochner-type inequalities. Recently, Petrotti [29] has analyzed how contractive coupling rates can be used to prove stronger inequalities in the form of curvature lower bounds for Markov chains and geodesic convexity of entropy functionals.
Maas and Matthes [25] suggested a finite-volume discretization for a nonlinear parabolic fourth-order equation, which can be written as a gradient flow similarly as the Fokker–Planck equation (1). The long-time asymptotics was shown for additive potentials by taking advantage of the factorization of the problem. Cancès and Venel [10] studied an implicit Euler finite-volume approximations of nonlinear Fokker–Planck equations using a discretization of the flux, which was motivated by probabilistic exclusion processes (and which have some similarity to our rates (5)). Their approach allows for non-additive potentials, but the long-time asymptotics was not investigated. The finite-volume scheme of [13] allows for the proof of exponential equilibration, but the decay rate depends on the size of the domain (because of the use of the Poincaré inequality; also see [19]), while we assume strong convexity of the potential. A link between finite-volume-type discretizations for the Fokker–Planck equation in Voronoï meshes and gradient flows with discrete Wasserstein distances on graphs was established in [1].
@@ 2. Definition of the continuous-time Markov chain
with center $i=(i_{1},\ldots,i_{d})\in{\mathcal{K}}_{h}$, where $i_{j}=(-K+n_{j}+h/2)$ for $n_{j}\in\{1,\ldots,2K/h\}$. Let $\sigma>0$ be a diffusion constant and $V\in C^{2}({\mathbb{R}}^{d})$ be a potential.
To define the continuous-time Markov chain, we recall definition (5) of the transition rates. The rate $c(i,\pm j)$ defines the numerical flux from the cell $N_{i}$ to $N_{i\pm he_{j}}$. The definition of $c(i,\pm_{j})$ implies that the numerical flux at the boundary vanishes. The rates are path-independent in the sense
@@
With these preparations, we can define the continuous-time Markov chain $(Y_{t}^{h})_{t\geq 0}$ on ${\mathcal{K}}_{h}$ via the discrete generator ${\mathcal{L}}_{h}$. Recall that $G_{i}$ denotes the set of all moves from $i$ to $i\pm he_{j}$ for $j=1,\ldots,d$, written as ±$\pm_{j}$. Let $(Z_{n}^{h})_{n\in{\mathbb{N}}}$ be a discrete-time Markov chain on ${\mathcal{K}}_{h}$ with initial distribtion $\nu$, given by the transition kernel
@@
The unique invariant measure $m_{h}$ of the continuous-time-time Markov chain $(Y_{t}^{h})_{t\geq 0}$ with the transition rates (5) is given by
@@
As the state space is finite and the Markov chain is irreducible and aperiodic, there exists a unique invariant measure $m_{h}$. To show that (14) is that invariant measure, we verify that $\sum_{i\in{\mathcal{K}}_{h}}{\mathcal{L}}_{h}f(i)m_{h}(i)=0$ for all positive functions $f$ on ${\mathcal{K}}_{h}$. We use summation by parts to factor out $f(i)$ and then insert definition (14):
$\sum_{i\in{\mathcal{K}}_{h}}$ ${\mathcal{L}}_{h}f(i)m_{h}(i)$
$=\sum_{i\in{\mathcal{K}}_{h}}\sum_{j=1}^{d}\big{[}c(i,+_{j})\big{(}f(i+he_{j})-f(i)\big{)}+c(i,-_{j})\big{(}f(i-he_{h})-f(i)\big{)}\big{]}m_{h}(i)$
$=\sum_{i\in{\mathcal{K}}_{h}}\sum_{j=1}^{d}f(i)\big{[}c(i-he_{j},+_{j})m_{h}(i-he_{j})+c(i+he_{j},-_{j})m_{h}(i+he_{j})$
$\phantom{xx}-c(i,+_{j})m_{h}(i)-c(i,-_{j})m_{h}(i)\big{]}$
$=\sum_{i\in{\mathcal{K}}_{h}}\sum_{j=1}^{d}\frac{f(i)}{Z}\Big{(}c(i-he_{j},+_{j})e^{-V^{h}(i-he_{j})/\sigma^{2}}\mathrm{1}_{\{i-he_{j}\in{\mathcal{K}}_{h}\}}-c(i,+_{j})e^{-V^{h}(i)/\sigma^{2}}$
$\phantom{xx}+c(i+he_{j},-_{j})e^{-V^{h}(i+he_{j})/\sigma^{2}}\mathrm{1}_{\{i+he_{j}\in{\mathcal{K}}_{h}\}}-c(i,-_{j})e^{-V^{h}(i)/\sigma^{2}}\Big{)}.$
Finally, we insert definition (5) of the rates $c$ and rearrange the terms:
$\sum_{i\in{\mathcal{K}}_{h}}$ ${\mathcal{L}}_{h}f(i)m_{h}(i)=\sum_{i\in{\mathcal{K}}_{h}}\sum_{j=1}^{d}\frac{\sigma^{2}}{h^{2}Z}f(i)\Big{(}\big{[}e^{-(V^{h}(i)-V^{h}(i-he_{j}))/(2\sigma^{2})}e^{-V^{h}(i-he_{j})/\sigma^{2}}$
$-e^{-(V^{h}(i-he_{j})-V^{h}(i))/(2\sigma^{2})}e^{-V^{h}(i)/\sigma^{2}}\big{]}\mathrm{1}_{\{i-he_{j}\in{\mathcal{K}}_{h}\}}$
$+\big{[}e^{-(V^{h}(i)-V^{h}(i+he_{j}))/(2\sigma^{2})}e^{-V^{h}(i+he_{j})/\sigma^{2}}$
$-e^{-(V^{h}(i+he_{j})-V^{h}(i))/(2\sigma^{2})}e^{-V^{h}(i)/\sigma^{2}}\big{]}\mathrm{1}_{\{i+he_{j}\in{\mathcal{K}}_{h}\}}\Big{)}=0,$
since the expressions in the brackets []delimited-[][$\cdots$][ ⋯ ] vanish. ∎
We remark that the continuous-time Markov chain $(Y_{n}^{h})_{n\in{\mathbb{N}}}$ satisfies the detailed-balance condition $c(i,+_{j})m_{h}(i)=c(i+he_{j},-_{j})m_{h}(i+he_{j})$ for all $i\in{\mathcal{K}}_{h}$ and $j=1,\ldots,d$, which implies the reversibility of the continuous-time Markov chain starting with initial measure $m_{h}$.
@@
where $\mbox{where }M_{h}(i,\pm_{j})=\sqrt{m_{h}(i\pm he_{j})m_{h}(i)}.$
The mean value $M_{h}(i,\pm_{j})$ corresponds to the geometric mean used in [26, Cor. 5.5] (also see [23, Sec. 4]). Other choices for $M_{h}(i,\pm_{j})$ can be found in [20, Sec. 3.2] and [13, Sec. 2.1.3]. We observe that the difference of the flows to and from the neighboring cells satisfies
@@ 3. Main results for continuous-time Markov chains
We present the main results for continuous-time Markov chains. The proofs will be given in Appendix A and Section 6.
@@ 3.1. Convergence of the Markov chain approximation to the SDE
To construct a continuous-time Markov chain that converges to the solution to the SDE ${\mathrm{d}}X_{t}=-\nabla V(X_{t}){\mathrm{d}}t+\sqrt{2\sigma}{\mathrm{d}}B_{t}$ in the whole space, with $(B_{t})_{t\geq 0}$ being a Brownian motion, we need to modify the cube $D=[-K,K]^{d}$. In fact, we replace $K$ by $K_{h}$ for $h>0$ such that $K_{h}\to\infty$ as $h\to 0$. Then Theorem 3 holds similarly after adapting the diffusion approximation of [17, Theorem 7.4.1].
@@ 3.2. Exponential decay in $\phi$-entropy
Consider the Markov chain $(Y_{t}^{h})_{t\geq 0}$ on ${\mathcal{K}}_{h}$, characterized by the generator (10). Let ϕ:++:italic-$\phi:{\mathbb{R}}_{+}\to{\mathbb{R}}_{+}$ and $f,g:{\mathcal{K}}_{h}\to{\mathbb{R}}_{+}$ be functions, where ${\mathbb{R}}_{+}:=[0,\infty)$. We recall the definition of the $\phi$-entropy, relative to the invariant measure $m_{h}$:
@@
where we have set $\nabla_{\gamma}f(i)=f(\gamma i)-f(i)$. The last identity follows from the reversibility of the Markov chain (see, e.g., [15, (2.12)]).
@@
Then the convex Sobolev inequality becomes the modified logarithmic Sobolev inequality for $\alpha=1$, the Poincaré inequality for $\alpha=2$, and the Beckner inequality for $21<\alpha<$ 2.
@@
[…]
@@
Let Assumptions (A1)–(A3) hold. Consider the continuous-time Markov chain $(Y_{t}^{h})_{t\geq 0}$ on ${\mathcal{K}}_{h}$ with transition rates (5), which satisfy (9). Then the convex Sobolev inequality (8) holds with constant
$\kappa_{\phi}=2\min_{i,i+he_{j}\in{\mathcal{K}}_{h},\,j=1,\ldots,d}\{\kappa_{+}(i,j),\kappa_{-}(i+he_{j},j)\},$
@@
×$\phantom{xx}\times\bigg{[}1-\exp\bigg{(}-\frac{1}{2\sigma^{2}}\int_{0}^{h}\big{(}\partial V_{j}^{h}(i_{j}+h+s)-\partial V_{j}^{h}(i_{j}+s)\big{)}{\mathrm{d}}s\bigg{)}\bigg{]}.$
@@
Let ±$\partial_{\pm j}^{h}$ be the discrete derivative on ${\mathcal{K}}_{h}$ in direction $j$, defined by ±$\partial_{\pm j}^{h}f(i)=\pm(f(i+he_{j})-f(i))/h$. A Taylor expansion yields

### v2

@@ Abstract.
The authors acknowledge partial support from the Austrian Science Fund (FWF), grant 10.55776/F65, and from the Austrian Federal Ministry for Women, Science and Research and implemented by ÖAD, project MultHeFlo. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme, ERC Advanced Grant NEUROMORPH, no. 101018153. For open-access purposes, the authors have applied a $C$ BY public copyright license to any author-accepted manuscript version arising from this submission.
@@ 1.1. The setting
with the initial condition $X_{0}$ with law $u_{0}$, $e_{j}$ are the Euclidean basis vectors, and the local time $L^{j\pm}$ is a continuous, nondecreasing process such that $L_{0}^{j\pm}=0$, which increases only when the process $(X_{t})_{t\geq 0}$ is in the half-spaces associated to $D$, i.e. $L_{t}^{j\pm}=\int_{0}^{t}\mathrm{1}_{\{X_{s}\in\partial D_{j}^{\pm}\}}{\mathrm{d}}L_{s}^{j\pm}$, where $\partial D_{j}^{\pm}=\{x\in D:\langle x\mp Ke_{j},e_{j}\rangle=0\}$ and $\langle\cdot,\cdot\rangle$ is the standard inner product on ${\mathbb{R}}^{d}$. The existence and uniqueness of a strong solution to (2) holds by [32]. The terms involving $L^{j\pm}_{t}$ ensure that the solution to the SDE is reflected at the boundary $\partial D$.
@@
where $G_{i}$ is the set of all moves $\pm_{j}$ to the neighboring states $i\mapsto i\pm he_{j}$ (with the grid size $h>0$) and $c(i,\gamma)$ denotes the rate that a jump happens from $i$ in the direction $\gamma$,
@@ 1.2. Main results and key ideas
Condition (i) is satisfied for the power-law entropies $\phi_{\alpha}(x)=x^{\alpha}$ with $1<\alpha\leq 2$ and the Boltzmann entropy density $\phi_{\alpha}(x)=x\log x$ with $\alpha=1$. For this family of functions, the exponential decay is closely related to convex Sobolev inequalities, namely the modified logarithmic Sobolev inequality for $\alpha=1$, the Poincaré inequality for $\alpha=2$, and Beckner-type inequalities for $1<\alpha<2$. The convexity of $\phi$ is natural in this context, while the convexity of $\Phi$ is needed in Conforti’s coupling approach; see Theorem 16 below. The strong convexity of the potential in condition (ii) is required to conclude exponential decay. Condition (iii) is satisfied, for instance, for additive potentials being of the form $V(x)=\sum_{j=1}^{d}V_{j}(x_{j})$ with $x=(x_{1},\ldots,x_{d})\in{\mathbb{R}}^{d}$; see Remark 5. An example of a non-additive potential satisfying condition (iii) is given by $V(x)=x^{T}Kx$, where $K$ is a symmetric positive definite matrix with smallest eigenvalue $\kappa>0$.
Our main results include the following quantitative decay rates for the continuous-time Markov chain and its discrete-time version (obtained from an explicit Euler scheme) in the relative $\phi$-entropy
@@
(6) $\mathcal{H}^{\phi}(S_{t}f|m_{h})\leq e^{-\kappa_{\phi}t}\mathcal{H}^{\phi}(f|m_{h})$ $\quad\mbox{for }f\geq 0,\ t\geq 0,$
(7) $\mathcal{W}_{1}(\nu p_{t},\eta p_{t})\leq\sqrt{d}e^{-\kappa_{1}t}\mathcal{W}_{1}(\nu,\eta)$ $\quad\mbox{for }t\geq 0,$
@@
Therefore, the second idea is to use the coupling method of Conforti [15]. Roughly speaking, the idea is to compare two probability measures $\nu$ and $\eta$ and to construct a joint probability spaces with marginals $\nu$ and $\eta$. Conforti proved (6) for various interacting random walks. We establish (6) in the context of finite-volume schemes with strongly convex potentials and interpret the value $\kappa_{\phi}$ within this setting. In the proof, we use a synchronous coupling for the contraction rates with maximal probability. When a synchronous move is not possible, we couple the rates of the neighboring states such that the states after a jump are identical with maximal probability or are at least neighboring states.
Inequality (7) is proved by using the path coupling method, see [9]. This coupling approach reduces the computations in the proof only to neighbouring states. It holds for two continuous-time Markov chains $Y_{t}^{1}$ and $Y_{t}^{2}$ that
@@
We prove similar results as (6)–(7) for discrete-time Markov chains associated to an explicit Euler scheme for the Fokker–Planck equation. Here, the idea is to apply the discrete Bakry–Emery method of [23], which is based on an estimate between the Fisher information $\mathcal{E}(\phi^{\prime}(f),f)$ and the entropy production $({\mathrm{d}}\mathcal{H}^{\phi}/{\mathrm{d}}t)(S_{t}f|m_{h})$ (which are the same in the continuous setting). We do not obtain contraction but exponential decay with a prefactor that depends on the initial datum $f$.
@@ 1.3. Comparison to the literature
Mielke investigated in [27] geodesic convexity properties of nonlocal transportation distances on probability spaces such that continuous-time Markov chains can be formulated as gradient flows. It is known that geodesic convexity implies exponential decay [2]. This idea was extended to nonlinear Fokker–Planck equations in [13].
@@
Conforti established in [15] a new probabilistic approach to prove convex Sobolev inequalities of the type (8) and to quantify the exponential decay in terms of the entropy for continuous-time Markov chains. His approach uses coupling rates to treat the second time derivative of the entropy, thus avoiding the use of discrete Bochner-type inequalities. Recently, Pedrotti [31] has analyzed how contractive coupling rates can be used to prove stronger inequalities in the form of curvature lower bounds for Markov chains and geodesic convexity of entropy functionals.
Maas and Matthes [26] suggested a finite-volume discretization for a nonlinear parabolic fourth-order equation, which can be written as a gradient flow similarly as the Fokker–Planck equation (1). The long-time asymptotics was shown for additive potentials by taking advantage of the factorization of the problem. Cancès and Venel [10] studied an implicit Euler finite-volume approximation of nonlinear Fokker–Planck equations using a discretization of the flux, which was motivated by probabilistic exclusion processes (and which have some similarity to our rates (5)). Their approach allows for non-additive potentials, but the long-time asymptotics was not investigated. The finite-volume scheme of [14] allows for the proof of exponential equilibration, but the decay rate depends on the size of the domain (because of the use of the Poincaré inequality; also see [20]), while we assume strong convexity of the potential. A link between finite-volume-type discretizations for the Fokker–Planck equation in Voronoï meshes and gradient flows with discrete Wasserstein distances on graphs was established in [1].
@@ 2. Definition of the continuous-time Markov chain
with center $i=(i_{1},\ldots,i_{d})\in{\mathcal{K}}_{h}$, where $i_{j}=-K+hn_{j}-h/2$ for $n_{j}\in\{1,\ldots,2K/h\}$. Let $\sigma>0$ be a diffusion constant and $V\in C^{2}({\mathbb{R}}^{d})$ be a potential.
To define the continuous-time Markov chain, we recall definition (5) of the transition rates. The rate $c(i,\pm_{j})$ defines the numerical flux from the cell $N_{i}$ to $N_{i\pm he_{j}}$. The definition of $c(i,\pm_{j})$ implies that the numerical flux at the boundary vanishes. The rates are path-independent in the sense
@@
With these preparations, we can define the continuous-time Markov chain $(Y_{t}^{h})_{t\geq 0}$ on ${\mathcal{K}}_{h}$ via the discrete generator ${\mathcal{L}}_{h}$. Recall that $G_{i}$ denotes the set of all moves from $i$ to $i\pm he_{j}$ for $j=1,\ldots,d$, written as $\pm_{j}$. Let $(Z_{n}^{h})_{n\in{\mathbb{N}}}$ be a discrete-time Markov chain on ${\mathcal{K}}_{h}$ with initial distribtion $\nu$, given by the transition kernel
@@
The unique invariant measure $m_{h}$ of the continuous-time Markov chain $(Y_{t}^{h})_{t\geq 0}$ with the transition rates (5) is given by
@@
The result follows immediately from the fact that the continuous-time Markov chain $(Y_{n}^{h})_{n\in{\mathbb{N}}}$ satisfies the detailed-balance condition $c(i,+_{j})m_{h}(i)=c(i+he_{j},-_{j})m_{h}(i+he_{j})$ for all $i\in{\mathcal{K}}_{h}$ and $j=1,\ldots,d$, which implies both the reversibility of the continuous-time Markov chain starting with initial measure $m_{h}$ and that $m_{h}$ given in (14) is the unique invariant measure of $(Y_{t}^{h})_{t\geq 0}$.
Alternatively, since the state space is finite and the Markov chain is irreducible and aperiodic, there exists a unique invariant measure $m_{h}$ and we can verify that $\sum_{i\in{\mathcal{K}}_{h}}{\mathcal{L}}_{h}f(i)m_{h}(i)=0$ holds for all positive functions $f$ on ${\mathcal{K}}_{h}$, which shows that (14) is the invariant measure. ∎
@@
$\mbox{where }M_{h}(i,\pm_{j})=\sqrt{m_{h}(i\pm he_{j})m_{h}(i)}.$
The mean value $M_{h}(i,\pm_{j})$ corresponds to the geometric mean used in [27, Cor. 5.5] (also see [24, Sec. 4]). Other choices for $M_{h}(i,\pm_{j})$ can be found in [21, Sec. 3.2] and [14, Sec. 2.1.3]. We observe that the difference of the flows to and from the neighboring cells satisfies
@@ 3. Main results for continuous-time Markov chains
We present the main results for continuous-time Markov chains. The proofs will be given in Section 6 and Appendix A.
@@ 3.1. Convergence of the Markov chain approximation to the SDE
To construct a continuous-time Markov chain that converges to the solution to the SDE ${\mathrm{d}}X_{t}=-\nabla V(X_{t}){\mathrm{d}}t+\sqrt{2\sigma}{\mathrm{d}}B_{t}$ in the whole space, with $(B_{t})_{t\geq 0}$ being a Brownian motion, we need to modify the cube $D=[-K,K]^{d}$. In fact, we replace $K$ by $K_{h}$ for $h>0$ such that $K_{h}\to\infty$ as $h\to 0$. Then Theorem 3 holds similarly after adapting the diffusion approximation of [18, Theorem 7.4.1].
@@ 3.2. Exponential decay in $\phi$-entropy
Consider the Markov chain $(Y_{t}^{h})_{t\geq 0}$ on ${\mathcal{K}}_{h}$, characterized by the generator (10). Let $\phi:{\mathbb{R}}_{+}\to{\mathbb{R}}_{+}$ and $f,g:{\mathcal{K}}_{h}\to{\mathbb{R}}_{+}$ be functions, where ${\mathbb{R}}_{+}:=[0,\infty)$. We recall the definition of the $\phi$-entropy, relative to the invariant measure $m_{h}$:
@@
where we have set $\nabla_{\gamma}f(i)=f(\gamma i)-f(i)$. The last identity follows from the reversibility of the Markov chain (see, e.g., [16, (2.12)]).
@@
Then the convex Sobolev inequality becomes the modified logarithmic Sobolev inequality for $\alpha=1$, the Poincaré inequality for $\alpha=2$, and the Beckner inequality for $1<\alpha<2$.
@@
    Strong $\kappa$-convexity: Let $V\in C^{1}(D;{\mathbb{R}})$ be such that there exists $\kappa>0$ such that $\langle x-y,\nabla V(x)-\nabla V(y)\rangle\geq\kappa|x-y|^{2}$ for all $x,y\in D$, where $\langle\cdot,\cdot\rangle$ is the inner product on ${\mathbb{R}}^{d}$.
@@
Let Assumptions (A1) and (A3) hold. Consider the continuous-time Markov chain $(Y_{t}^{h})_{t\geq 0}$ on ${\mathcal{K}}_{h}$ with transition rates (5), which satisfy (9). Then the convex Sobolev inequality (8) holds with constant
$\kappa_{\phi}=\min_{i,i+he_{j}\in{\mathcal{K}}_{h},\,j=1,\ldots,d}\{\kappa_{+}(i,j)+\kappa_{-}(i+he_{j},j)\},$
@@
$\phantom{xx}\times\bigg[1-\exp\bigg(-\frac{1}{2\sigma^{2}}\int_{0}^{h}\big(\partial V_{j}^{h}(i_{j}+h+s)-\partial V_{j}^{h}(i_{j}+s)\big){\mathrm{d}}s\bigg)\bigg].$
@@
Let $\partial_{\pm j}^{h}$ be the discrete derivative on ${\mathcal{K}}_{h}$ in direction $j$, defined by $\partial_{\pm j}^{h}f(i)=\pm(f(i+he_{j})-f(i))/h$. A Taylor expansion yields